# appendixtocaptAppendix

### v3

@@ B.4 Connection to Geodesic Distance
While Varadhan’s original large deviation principle [60] applies to smooth manifolds without boundaries, extensions to domains with reflecting boundaries [43] ensure that the short-time behavior of the heat kernel $p(y,t|x,0)$ reflects the geodesic distance $d_{g}(x,y)$, the shortest path within $\Omega$ avoiding obstacles. For our normalized transition probability: a similar asymptotic relationship holds:
$\lim_{t\to 0}-t\log q(y|x,t)=\frac{d_{g}^{2}(x,y)}{4\alpha}$ (35)
where $\alpha=1/4$ is the diffusion coefficient. This follows because $p(x|x,t)$ and $p(y|y,t)$, influenced by boundary reflections, have asymptotic forms that cancel in the logarithm as $t\to 0$, leaving the geodesic term dominant.
As $t$ increases, our distance metric transitions from approximating geodesic distance to incorporating more global aspects of the domain’s connectivity, creating a multi-scale representation that seamlessly integrates local metric information with global connectivity structure. Section D explains eigen analysis of topological connectivity.
Appendix C Open-Field Environment
In unbounded, obstacle-free environments (open fields), the symmetric random walk simplifies to an isotropic diffusion process, offering a theoretical justification for the Gaussian tuning of place cells observed in such settings [45, 64]. As the grid discretization refines ($dx\to 0$, $dt\to 0$, $dx=\sqrt{dt}$), the transition probability $p(y|x,t)$ converges to the heat equation’s fundamental solution in 2D free space:
$p(y|x,t)=\frac{1}{4\pi\alpha t}\exp\left(-\frac{\|y-x\|^{2}}{4\alpha t}\right)$ (36)
where $\alpha=1/4$ is the diffusion coefficient, and $\|y-x\|^{2}$ is the squared Euclidean distance. Since $p(x|x,t)=\frac{1}{4\pi\alpha t}$ is position-independent, the normalized transition probability becomes:
$q(y|x,t)=\frac{p(y|x,t)}{\sqrt{p(x|x,t)p(y|y,t)}}=\exp\left(-\frac{\|y-x\|^{2}}{4\alpha t}\right)=\exp\left(-\frac{\|y-x\|^{2}}{t}\right)$ (37)
This Gaussian kernel, with variance $\sigma^{2}=2\alpha t=t/2$, reflects the diffusive spread of the random walk and mirrors the approximately Gaussian firing fields of hippocampal place cells in open environments. Given $q(y|x,t)=\langle h(x,t),h(y,t)\rangle$, the position embeddings $h(x,t)$ must reproduce this Gaussian decay. We construct $h(x,t)$ with Gaussian components, providing a mathematical basis for why place cell population activity exhibits Gaussian profiles in open fields.
@@ Appendix C Open-Field Environment
In an unbounded 2D open field, where $q(y|x,t)=\exp\left(-\frac{\|y-x\|^{2}}{t}\right)$, there exists a position embedding $h(x,t)\in\mathbb{R}^{n}$ with non-negative components such that $\langle h(x,t),h(y,t)\rangle=q(y|x,t)$, and each component $h_{i}(x,t)$ is a Gaussian function of $x$ with variance $t/2$ per dimension.
@@
The transition kernel $q(y|x,t)=\exp\left(-\frac{\|y-x\|^{2}}{t}\right)$ is a positive definite Gaussian kernel with variance $t/2$ (since $4\alpha t=t$ for $\alpha=1/4$), consistent with the Gaussian tuning of place cells in open fields. We construct $h(x,t)\in\mathbb{R}^{n}$ directly as:
$h_{i}(x,t)=c_{i}\exp\left(-\frac{\|x-\mu_{i}\|^{2}}{t}\right)$ (38)
where $\mu_{i}\in\mathbb{R}^{2}$ are fixed anchor points (e.g., a uniform grid), $c_{i}>0$ are constants, and the variance per dimension is $t/2$. The inner product is:
$\langle h(x,t),h(y,t)\rangle$ $=\sum_{i=1}^{n}h_{i}(x,t)h_{i}(y,t)=\sum_{i=1}^{n}c_{i}^{2}\exp\left(-\frac{\|x-\mu_{i}\|^{2}+\|y-\mu_{i}\|^{2}}{t}\right)$ (39)
@@
$\langle h(x,t),h(y,t)\rangle=\exp\left(-\frac{\|x-y\|^{2}}{t}\right)\sum_{i=1}^{n}c_{i}^{2}\exp\left(-\frac{2\left\|\frac{x+y}{2}-\mu_{i}\right\|^{2}}{t}\right)$ (41)
For a dense set of $\mu_{i}$ (e.g., a grid with spacing tmuch-less-$\ll\sqrt{t}$), the sum approximates a constant over a local region around $(x+y)/2$, as the terms $\exp\left(-\frac{2\left\|\frac{x+y}{2}-\mu_{i}\right\|^{2}}{t}\right)$ form a kernel density estimate. Set $c_{i}^{2}=\frac{1}{n}$; as $n\to\infty$, the sum converges to a constant $C$ (proportional to the density of $\mu_{i}$), yielding:
$\langle h(x,t),h(y,t)\rangle\approx C\exp\left(-\frac{\|x-y\|^{2}}{t}\right)$ (42)
Adjust $C=1$ by scaling $c_{i}$ (e.g., $c_{i}=\sqrt{\frac{1}{Cn}}$), ensuring $\langle h(x,t),h(y,t)\rangle=q(y|x,t)$. Each $h_{i}(x,t)$ is Gaussian with variance $t/2$ per dimension, and $h_{i}(x,t)\geq 0$, satisfying the theorem. ∎
@@
Appendix D Path Planning Properties
@@
D.1 Scale Transition Dynamics
The optimal scale $t^{*}=\operatorname*{arg\,max}_{t}\|\nabla_{x}q(x|y,t)\|$ adapts dynamically to the agent’s distance from the goal, ensuring efficient navigation. In an open field, we formalize this selection with the following theorem.
@@ D.1 Scale Transition Dynamics
In an open field with a Gaussian transition kernel $p(x|y,t)=\frac{1}{4\pi\alpha t}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha t}\right)$, where $\alpha=1/4$, the optimal time scale $t^{*}$ that maximizes the gradient magnitude $\|\nabla_{x}q(x|y,t)\|$, with $q(x|y,t)={p(x|y,t)}/{\sqrt{p(x|x,t)\cdot p(y|y,t)}}$, is:
$t^{*}=\frac{d^{2}(x,y)}{4\alpha}=d^{2}(x,y)$ (43)
@@
$p(x|y,t)=\frac{1}{4\pi\alpha t}\exp\left(-\frac{d^{2}}{4\alpha t}\right),\quad d=d(x,y)$ (44)
@@
$p(x|x,t)=p(y|y,t)=\frac{1}{4\pi\alpha t},\quad\sqrt{p(x|x,t)\cdot p(y|y,t)}=\frac{1}{4\pi\alpha t}$ (45)
@@
$q(x|y,t)=\frac{p(x|y,t)}{\frac{1}{4\pi\alpha t}}=\exp\left(-\frac{d^{2}}{4\alpha t}\right)$ (46)
@@
$\nabla_{x}q(x|y,t)=\exp\left(-\frac{d^{2}}{4\alpha t}\right)\cdot\left(-\frac{2(x-y)}{4\alpha t}\right)=-\frac{x-y}{2\alpha t}\exp\left(-\frac{d^{2}}{4\alpha t}\right)$ (47)
@@
$\|\nabla_{x}q(x|y,t)\|=\frac{d}{2\alpha t}\exp\left(-\frac{d^{2}}{4\alpha t}\right)$ (48)
Maximize $f(t)=\frac{d}{2\alpha t}\exp\left(-\frac{d^{2}}{4\alpha t}\right)$. Compute:
$\ln f(t)=\ln\left(\frac{d}{2\alpha}\right)-\ln t-\frac{d^{2}}{4\alpha t}$ (49)
@@
$\frac{\partial\ln f}{\partial t}=-\frac{1}{t}+\frac{d^{2}}{4\alpha t^{2}}=0\implies t=\frac{d^{2}}{4\alpha}$ (50)
For $\alpha=1/4$, $t^{*}=d^{2}$. The second derivative at $t^{*}$:
$\frac{\partial^{2}\ln f}{\partial t^{2}}=\frac{1}{t^{2}}-\frac{d^{2}}{2\alpha t^{3}},\quad\text{at }t=\frac{d^{2}}{4\alpha},\quad\frac{d^{2}}{2\alpha t}=2,\quad\frac{\partial^{2}\ln f}{\partial t^{2}}=-\frac{1}{t^{2}}<0$ (51)
@@
In an open field, $t^{*}\propto d^{2}(x,y)$, so $t^{*}$ decreases as the agent approaches the goal, focusing on finer spatial scales for precision.
D.2 Properties of the Gradient Field
The gradient field $\nabla_{x}q(x|y,t)$ drives navigation. We highlight three properties ensuring a smooth, trap-free path to the goal $y$.
First, $p(x|y,t)$ satisfies the heat equation with reflecting boundary conditions:
$\frac{\partial p(x|y,t)}{\partial t}=\alpha\nabla^{2}p(x|y,t),\quad\alpha=1/4,\quad\frac{\partial p(x|y,t)}{\partial n}=0\text{ on }\partial\Omega_{\text{obstacles}}$ (52)
For fixed $y$ and $t$, $p(x|y,t)$ is smooth in $x$, as the heat kernel (e.g., $p(x|y,t)=\frac{1}{4\pi\alpha t}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha t}\right)$ in an open field) is infinitely differentiable [24]. It has a unique maximum at $x=y$.
Second, $p(x|x,t)$ is smooth in $x$ for fixed $t$. In an open field, $p(x|x,t)=\frac{1}{4\pi\alpha t}$ is constant, while in general, $p(x|x,t)$ varies smoothly due to the heat kernel’s differentiability, reflecting the domain’s geometry near obstacles.
Third, since the random walk is symmetric, $\nabla_{x}q(x|y,t)=\nabla_{x}q(y|x,t)$. The gradient field of $q(y|x,t)$ is smooth, as $q(y|x,t)=q(x|y,t)$ inherits the smoothness of $p(x|y,t)$, and has a unique maximum at $x=y$.
These properties ensure that $\nabla_{x}q(y|x,t)$ forms a smooth field with a unique maximum of $q(y|x,t)$ at the goal $x=y$, producing a continuous, trap-free path toward $y$. This mirrors the hippocampus’s efficient navigation, where place cells encode smooth, goal-directed trajectories [38].
D.3 Planned Path vs. Shortest Path
The planned path follows the gradient $\nabla_{x}q(x|y,t)$, and $p(x|y,t)$ satisfies the heat equation with reflecting boundary conditions. We compare this path to the shortest (geodesic) path from $x$ to the goal $y$.

### v4

@@ D.4 Connection to Geodesic Distance
While Varadhan’s original large deviation principle [16] applies to smooth manifolds without boundaries, extensions to domains with reflecting boundaries [61] ensure that the short-time behavior of the heat kernel $p(y,\tau|x,0)$ reflects the geodesic distance $d_{g}(x,y)$, the shortest path within $\Omega$ avoiding obstacles. For our normalized transition probability: a similar asymptotic relationship holds:
$\lim_{\tau\to 0}-\tau\log q(y|x,\tau)=\frac{d_{g}^{2}(x,y)}{4\alpha},$ (32)
where $\alpha=1/4$ is the diffusion coefficient. This follows because $p(x|x,\tau)$ and $p(y|y,\tau)$, influenced by boundary reflections, have asymptotic forms that cancel in the logarithm as $\tau\to 0$, leaving the geodesic term dominant.
As $\tau$ increases, our distance metric transitions from approximating geodesic distance to incorporating more global aspects of the domain’s connectivity, creating a multi-scale representation that seamlessly integrates local metric information with global connectivity structure. Section G explains eigen analysis of connectivity.
Appendix E Open-Field Environment
In unbounded, obstacle-free environments (open fields), the symmetric random walk simplifies to an isotropic diffusion process, offering a theoretical justification for the Gaussian tuning of place cells observed in such settings [1, 29]. As the grid discretization refines ($dx\to 0$, $dt\to 0$, $dx=\sqrt{dt}$), the transition probability $p(y|x,\tau)$ converges to the heat equation’s fundamental solution in 2D free space:
$p(y|x,\tau)=\frac{1}{4\pi\alpha\tau}\exp\left(-\frac{\|y-x\|^{2}}{4\alpha\tau}\right)$ (33)
where $\alpha=1/4$ is the diffusion coefficient, and $\|y-x\|^{2}$ is the squared Euclidean distance. Since $p(x|x,\tau)=\frac{1}{4\pi\alpha\tau}$ is position-independent, the normalized transition probability becomes:
$q(y|x,\tau)=\frac{p(y|x,\tau)}{\sqrt{p(x|x,\tau)p(y|y,\tau)}}=\exp\left(-\frac{\|y-x\|^{2}}{4\alpha\tau}\right)=\exp\left(-\frac{\|y-x\|^{2}}{\tau}\right)$ (34)
This Gaussian kernel, with variance $\sigma^{2}=2\alpha\tau=\tau/2$, reflects the diffusive spread of the random walk and mirrors the approximately Gaussian firing fields of hippocampal place cells in open environments. Given $q(y|x,\tau)=\langle h(x,\tau),h(y,\tau)\rangle$, the position embeddings $h(x,\tau)$ must reproduce this Gaussian decay. We construct $h(x,\tau)$ with Gaussian components, providing a mathematical basis for why place cell population activity exhibits Gaussian profiles in open fields.
@@ Appendix E Open-Field Environment
In an unbounded 2D open field, where $q(y|x,\tau)=\exp\left(-\frac{\|y-x\|^{2}}{\tau}\right)$, there exists a position embedding $h(x,\tau)\in\mathbb{R}^{n}$ with non-negative components such that $\langle h(x,\tau),h(y,\tau)\rangle=q(y|x,\tau)$, and each component $h_{i}(x,\tau)$ is a Gaussian function of $x$ with variance $\tau/2$ per dimension.
@@
The transition kernel $q(y|x,\tau)=\exp\left(-\frac{\|y-x\|^{2}}{\tau}\right)$ is a positive definite Gaussian kernel with variance $\tau/2$ (since $4\alpha\tau=\tau$ for $\alpha=1/4$), consistent with the Gaussian tuning of place cells in open fields. We construct $h(x,\tau)\in\mathbb{R}^{n}$ directly as:
$h_{i}(x,\tau)=c_{i}\exp\left(-\frac{\|x-\mu_{i}\|^{2}}{\tau}\right)$ (35)
where $\mu_{i}\in\mathbb{R}^{2}$ are fixed anchor points (e.g., a uniform grid), $c_{i}>0$ are constants, and the variance per dimension is $\tau/2$. The inner product is:
$\langle h(x,\tau),h(y,\tau)\rangle$ $=\sum_{i=1}^{n}h_{i}(x,\tau)h_{i}(y,\tau)=\sum_{i=1}^{n}c_{i}^{2}\exp\left(-\frac{\|x-\mu_{i}\|^{2}+\|y-\mu_{i}\|^{2}}{\tau}\right)$ (36)
@@
$\langle h(x,\tau),h(y,\tau)\rangle=\exp\left(-\frac{\|x-y\|^{2}}{\tau}\right)\sum_{i=1}^{n}c_{i}^{2}\exp\left(-\frac{2\left\|\frac{x+y}{2}-\mu_{i}\right\|^{2}}{\tau}\right)$ (38)
For a dense set of $\mu_{i}$ (e.g., a grid with spacing $\ll\sqrt{\tau}$), the sum approximates a constant over a local region around $(x+y)/2$, as the terms $\exp\left(-\frac{2\left\|\frac{x+y}{2}-\mu_{i}\right\|^{2}}{\tau}\right)$ form a kernel density estimate. Set $c_{i}^{2}=\frac{1}{n}$; as $n\to\infty$, the sum converges to a constant $C$ (proportional to the density of $\mu_{i}$), yielding:
$\langle h(x,\tau),h(y,\tau)\rangle\approx C\exp\left(-\frac{\|x-y\|^{2}}{\tau}\right)$ (39)
Adjust $C=1$ by scaling $c_{i}$ (e.g., $c_{i}=\sqrt{\frac{1}{Cn}}$), ensuring $\langle h(x,\tau),h(y,\tau)\rangle=q(y|x,\tau)$. Each $h_{i}(x,\tau)$ is Gaussian with variance $\tau/2$ per dimension, and $h_{i}(x,\tau)\geq 0$, satisfying the theorem. ∎
@@
Appendix F Emergent Sparsity from Non-negativity and Orthogonality
Let $h(x,\tau)\in\mathbb{R}_{+}^{n}$ denote the non-negative representation of position $x\in\mathcal{X}$ at scale $\sqrt{\tau}$, and let
$\langle h(x,\tau),\,h(y,\tau)\rangle\;=\;q(y\mid x,\tau)$ (40)
where $q(\cdot\mid\cdot,\tau)\geq 0$ is a symmetric multi-step transition kernel (adjacency at scale $\sqrt{\tau}$). Define the (undirected) transition graph $\mathcal{G}_{\tau}=(\mathcal{X},E_{\tau})$ by
$(x,y)\in E_{\tau}\quad\Longleftrightarrow\quad q(y\mid x,\tau)>0$ (41)
Note that this transition graph is for $\tau$ steps instead of one step. Write $\mathrm{supp}(v)=\{i\in[n]:v_{i}>0\}$ for the support of a non-negative vector $v\in\mathbb{R}^{n}_{+}$, and put
[…]
Let $\deg_{\tau}(x)$ denote the degree of $x$ in $\mathcal{G}_{\tau}$, $\Delta(\tau)=\max_{x}\deg_{\tau}(x)$ the maximum degree, and
$\rho(\tau)\;=\;\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\deg_{\tau}(x)$ (43)
the mean neighborhood size.
Lemma 2 (Non-negativity +\,+ orthogonality $\Rightarrow$ disjoint support).
If $h(x,\tau),h(y,\tau)\in\mathbb{R}_{+}^{n}$ and $\langle h(x,\tau),h(y,\tau)\rangle=0$, then
$\mathrm{supp}\big(h(x,\tau)\big)\;\cap\;\mathrm{supp}\big(h(y,\tau)\big)\;=\;\varnothing$ (44)
Proof.
By non-negativity,
$0=\langle h(x,\tau),h(y,\tau)\rangle=\sum_{i=1}^{n}h_{i}(x,\tau)\,h_{i}(y,\tau)$ (45)
is a sum of non-negative terms. Hence each term vanishes: $h_{i}(x,\tau)\,h_{i}(y,\tau)=0$ for all $i\in[n]$, and thus no coordinate is simultaneously positive in both vectors. ∎
Proposition 3 (Clique structure and sparsity bound).
Assume (40) and that
$q(y\mid x,\tau)=0\quad\text{whenever}\quad(x,y)\notin E_{\tau}$ (46)
Then for each coordinate $i\in[n]$:
  1. (i)
    $S_{i}(\tau)$ induces a clique in $\mathcal{G}_{\tau}$; i.e., every pair $x,y\in S_{i}(\tau)$ is adjacent in $\mathcal{G}_{\tau}$
  2. (ii)
    Consequently,
    $|S_{i}(\tau)|\;\leq\;1+\min_{x\in S_{i}(\tau)}\deg_{\tau}(x)\;\leq\;1+\Delta(\tau)$ (47)
  3. (iii)
    The average number of active coordinates per position obeys
    $\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\|h(x,\tau)\|_{0}\;=\;\frac{1}{|\mathcal{X}|}\sum_{i=1}^{n}|S_{i}(\tau)|\;\leq\;\frac{n\,(1+\Delta(\tau))}{|\mathcal{X}|}$ (48)
    Moreover, since $\Delta(\tau)=\max_{x}\deg_{\tau}(x)$ and $\rho(\tau)$ is the mean degree, any a priori bound of the form $\Delta(\tau)\leq C\,\rho(\tau)$ (with graph-dependent constant $C\geq 1$) yields
    $\frac{1}{|\mathcal{X}|}\sum_{x}\|h(x,\tau)\|_{0}\;\leq\;\frac{n\,[1+C\,\rho(\tau)]}{|\mathcal{X}|}$ (49)
Proof.
(i) Take any $x,y\in S_{i}(\tau)$ with $x\neq y$. Then $h_{i}(x,\tau)>0$ and $h_{i}(y,\tau)>0$, so by Lemma 2 they cannot be orthogonal. By the hypothesis $q(y\mid x,\tau)=\langle h(x,\tau),h(y,\tau)\rangle$, orthogonality occurs exactly when $(x,y)\notin E_{\tau}$. Therefore $(x,y)\in E_{\tau}$. Since the choice of $x,y$ was arbitrary, $S_{i}(\tau)$ induces a clique.
(ii) Fix any $x^{\ast}\in S_{i}(\tau)$. By (i), every $y\in S_{i}(\tau)\setminus\{x^{\ast}\}$ must be adjacent to $x^{\ast}$, so $S_{i}(\tau)\subseteq\{x^{\ast}\}\cup\mathcal{N}_{\tau}(x^{\ast})$, where $\mathcal{N}_{\tau}(x^{\ast})$ is the neighborhood of $x^{\ast}$. Thus $|S_{i}(\tau)|\leq 1+\deg_{\tau}(x^{\ast})$. Minimizing over $x^{\ast}\in S_{i}(\tau)$ gives the first inequality; the second follows from $\deg_{\tau}(x^{\ast})\leq\Delta(\tau)$.
(iii) The identity
$\sum_{x\in\mathcal{X}}\|h(x,\tau)\|_{0}\;=\;\sum_{i=1}^{n}|S_{i}(\tau)|$ (50)
is a double-counting equality (both sides count the number of pairs $(x,i)$ with $h_{i}(x,\tau)>0$). Using (ii) and summing over $i$ yields $\sum_{i}|S_{i}(\tau)|\leq n\,[1+\Delta(\tau)]$, and dividing by $|\mathcal{X}|$ establishes (48). If an a priori comparison $\Delta(\tau)\leq C\,\rho(\tau)$ holds for the graph family under consideration, the final bound follows immediately. ∎
Geometric specialization.
If $\mathcal{X}\subset\mathbb{R}^{m}$ is $\eta$-separated (minimum inter-point distance $\geq\eta>0$) and $\mathcal{G}_{\tau}$ is a geometric threshold graph
$(x,y)\in E_{\tau}\quad\Longleftrightarrow\quad\|x-y\|\leq\sqrt{\tau}$ (51)
then each closed neighborhood fits inside a ball $B(x,\sqrt{\tau})$ of radius $\sqrt{\tau}$. A standard packing argument implies
$\deg_{\tau}(x)\;+\;1\;\leq\;C_{m}\,\Big(\tfrac{\sqrt{\tau}}{\eta}\Big)^{m}$ (52)
for a constant $C_{m}$ depending only on the ambient dimension $m$. Consequently,
$|S_{i}(\tau)|\;\leq\;C_{m}\,\Big(\tfrac{\sqrt{\tau}}{\eta}\Big)^{m}\qquad\frac{1}{|\mathcal{X}|}\sum_{x}\|h(x,\tau)\|_{0},\;\leq\;\frac{d\,C_{m}}{|\mathcal{X}|}\,\Big(\tfrac{\sqrt{\tau}}{\eta}\Big)^{m}$ (53)
Thus, as $\tau$ increases, the allowable clique size (hence the bound on average support) grows polynomially with the geometric volume of the $\sqrt{\tau}$-ball, quantitatively linking scale to reduced sparsity (broader fields).
Corollary 4 (Scale-dependent localization).
Assume $\tau\mapsto\mathcal{G}_{\tau}$ is monotone in the sense that $\tau_{1}<\tau_{2}$ implies $E_{\tau_{1}}\subseteq E_{\tau_{2}}$. Then $\Delta(\tau)$ and $\rho(\tau)$ are non-decreasing in $\tau$, and the upper bounds in Proposition 3 are non-decreasing in $\tau$, i.e., expected sparsity decreases with scale, corresponding to larger place fields.
Proof.
Monotonicity $E_{\tau_{1}}\subseteq E_{\tau_{2}}$ implies $\deg_{\tau_{1}}(x)\leq\deg_{\tau_{2}}(x)$ for all $x$, hence $\Delta(\tau_{1})\leq\Delta(\tau_{2})$ and $\rho(\tau_{1})\leq\rho(\tau_{2})$. Apply Proposition 3 (iii). ∎
Interpretation.
Lemma 2 enforces disjoint supports for non-adjacent locations; Proposition 3 shows that each coordinate $i$ can only support a clique of mutually adjacent positions and bounds the size of that clique by local neighborhood size. In geometric environments, the bound scales with the volume of a $\sqrt{\tau}$-ball, making the $\tau$–field-size relationship explicit. Corollary 4 then formalizes the empirical trend: smaller $\tau$ $\Rightarrow$ higher sparsity (smaller fields); larger $\tau$ $\Rightarrow$ lower sparsity (larger fields).
Appendix G Path Planning Properties
@@
G.1 Scale Transition Dynamics
The optimal scale $\tau^{*}=\operatorname*{arg\,max}_{t}\|\nabla_{x}q(x|y,\tau)\|$ adapts dynamically to the agent’s distance from the goal, ensuring efficient navigation. In an open field, we formalize this selection with the following theorem.
@@ G.1 Scale Transition Dynamics
In an open field with a Gaussian transition kernel $p(x|y,\tau)=\frac{1}{4\pi\alpha\tau}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha\tau}\right)$, where $\alpha=1/4$, the optimal time scale $\tau^{*}$ that maximizes the gradient magnitude $\|\nabla_{x}q(x|y,\tau)\|$, with $q(x|y,\tau)={p(x|y,\tau)}/{\sqrt{p(x|x,\tau)\cdot p(y|y,\tau)}}$, is:
$\tau^{*}=\frac{d^{2}(x,y)}{4\alpha}=d^{2}(x,y)$ (54)
@@
$p(x|y,\tau)=\frac{1}{4\pi\alpha\tau}\exp\left(-\frac{d^{2}}{4\alpha\tau}\right),\quad d=d(x,y)$ (55)
@@
$p(x|x,\tau)=p(y|y,\tau)=\frac{1}{4\pi\alpha\tau},\quad\sqrt{p(x|x,\tau)\cdot p(y|y,\tau)}=\frac{1}{4\pi\alpha\tau}$ (56)
@@
$q(x|y,\tau)=\frac{p(x|y,\tau)}{\frac{1}{4\pi\alpha\tau}}=\exp\left(-\frac{d^{2}}{4\alpha\tau}\right)$ (57)
@@
$\nabla_{x}q(x|y,\tau)=\exp\left(-\frac{d^{2}}{4\alpha\tau}\right)\cdot\left(-\frac{2(x-y)}{4\alpha\tau}\right)=-\frac{x-y}{2\alpha\tau}\exp\left(-\frac{d^{2}}{4\alpha\tau}\right)$ (58)
@@
$\|\nabla_{x}q(x|y,\tau)\|=\frac{d}{2\alpha\tau}\exp\left(-\frac{d^{2}}{4\alpha\tau}\right)$ (59)
Maximize $f(\tau)=\frac{d}{2\alpha\tau}\exp\left(-\frac{d^{2}}{4\alpha\tau}\right)$. Compute:
$\ln f(\tau)=\ln\left(\frac{d}{2\alpha}\right)-\ln\tau-\frac{d^{2}}{4\alpha\tau}$ (60)
@@
$\frac{\partial\ln f}{\partial\tau}=-\frac{1}{\tau}+\frac{d^{2}}{4\alpha\tau^{2}}=0\implies\tau=\frac{d^{2}}{4\alpha}$ (61)
For $\alpha=1/4$, $\tau^{*}=d^{2}$. The second derivative at $\tau^{*}$:
$\frac{\partial^{2}\ln f}{\partial\tau^{2}}=\frac{1}{\tau^{2}}-\frac{d^{2}}{2\alpha\tau^{3}},\quad\text{at }\tau=\frac{d^{2}}{4\alpha},\quad\frac{d^{2}}{2\alpha\tau}=2,\quad\frac{\partial^{2}\ln f}{\partial\tau^{2}}=-\frac{1}{\tau^{2}}<0$ (62)
@@
In an open field, $\tau^{*}\propto d^{2}(x,y)$, so $\tau^{*}$ decreases as the agent approaches the goal, focusing on finer spatial scales for precision.
G.2 Properties of the Gradient Field
The gradient field $\nabla_{x}q(x|y,\tau)$ drives navigation. We highlight three properties ensuring a smooth, trap-free path to the goal $y$.
First, $p(x|y,\tau)$ satisfies the heat equation with reflecting boundary conditions:
$\frac{\partial p(x|y,\tau)}{\partial\tau}=\alpha\nabla^{2}p(x|y,\tau),\quad\alpha=1/4,\quad\frac{\partial p(x|y,\tau)}{\partial n}=0\text{ on }\partial\Omega_{\text{obstacles}}.$ (63)
For fixed $y$ and $\tau$, $p(x|y,\tau)$ is smooth in $x$, as the heat kernel (e.g., $p(x|y,\tau)=\frac{1}{4\pi\alpha\tau}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha\tau}\right)$ in an open field) is infinitely differentiable [15]. It has a unique maximum at $x=y$.
Second, $p(x|x,\tau)$ is smooth in $x$ for fixed $\tau$. In an open field, $p(x|x,\tau)=\frac{1}{4\pi\alpha\tau}$ is constant, while in general, $p(x|x,\tau)$ varies smoothly due to the heat kernel’s differentiability, reflecting the domain’s geometry near obstacles.
Third, since the random walk is symmetric, $\nabla_{x}q(x|y,\tau)=\nabla_{x}q(y|x,\tau)$. The gradient field of $q(y|x,\tau)$ is smooth, as $q(y|x,\tau)=q(x|y,\tau)$ inherits the smoothness of $p(x|y,\tau)$, and has a unique maximum at $x=y$.
These properties ensure that $\nabla_{x}q(y|x,\tau)$ forms a smooth field with a unique maximum at the goal, producing a continuous, trap-free path toward $y$. This mirrors the hippocampus’s efficient navigation, where place cells encode smooth, goal-directed trajectories [11].
G.3 Planned Path vs. Shortest Path
The planned path follows the gradient $\nabla_{x}q(x|y,\tau)$, and $p(x|y,\tau)$ satisfies the heat equation with reflecting boundary conditions. We compare this path to the shortest (geodesic) path from $x$ to the goal $y$.